OpenStudy (anonymous):

Help! Suppose that (f(x+h)-f(x))/h = (-2h(x+4)-h^2)/h(x+h+4)^2(x+4)^2 Find the slope m of the tangent line at x+1. m=?

OpenStudy (anonymous):

at \(x=1\) ?

OpenStudy (anonymous):

Yes sorry I typed it wrong

OpenStudy (anonymous):

replace \(x\) by \(1\) and compute the limit

OpenStudy (anonymous):

is it \[\frac{-2h(x+4)-h^2}{h(x+h+4)^2(x+4)^2} \]??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then start with \[\frac{-2h(1+4)-h^2}{h(1+h+4)^2(1+4)^2}\]

OpenStudy (anonymous):

\[\frac{-10h-h^2}{h(h+5)^2\times 25}\]

OpenStudy (anonymous):

I got that. Then I simplified the bottom.

OpenStudy (anonymous):

\[\frac{h(-10-h)}{25h(h+5)^2}=\frac{-10-h}{25(h+5)^2}\] now replace \(h\) by zero to get your answer

OpenStudy (anonymous):

don't multiply out, that is a bad idea factor and cancel the \(h\) so you can replace \(h\) by \(0\) without getting a zero in the denominator

OpenStudy (anonymous):

O ok.

OpenStudy (anonymous):

I got m= -10/625